# Cell as Point: One-Stage Framework for Efficient Cell Tracking

### v1

@@ 4 Experiments and Results
Table 1: Comparison Experiments with Baselines. $\mathrm{ES}$, $\mathrm{EA}$, $\mathrm{ED}$, and $\mathrm{ESM}$ are abbreviations of $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, and $\mathrm{EdgeSemantic}$, respectively. Our framework CAP is TRA only, while previous frameworks GNN, EmbedTrack, and KIT-GE are SEG+TRA illustrated in Fig. 1 (a). Average inference time/seconds ($\mathrm{IT}$) and performance ($\mathrm{TRA}$) are visualized in Fig. 1 (b).
We conducted experiments to demonstrate the high efficacy and efficiency of our end-to-end, one-stage framework for cell tracking, CAP. While utilizing minimal inference time and without extra segmentation masks, the performance is competitive and even outperforms existing works. Also, we perform comprehensive ablation studies to verify the effectiveness of CAP components in detail.
@@ 4.1 Experimental Setup
We evaluate our proposed framework on DynamicNuclearNet Tracking in DeepCell [45] public database, HeLa, PC-3, 3T3, and RAW264 datasets divided by specimen, and ISBI cell tracking challenge [43, 62] public datasets, PhC-C2DH-U373 (U373), Fluo-N2DH-GOWT1 (GOWT1), and Fluo-C2DL-Huh7 (Huh7). These temporal data used for cell tracking, shot by different devices in various durations on diverse specimens, exhibit diverse morphologies and sequence characteristics. DeepCell contains $96$ sequences in total: $45$ for HeLa, $7$ for PC-3, $10$ for 3T3, and $34$ for RAW264, with $42$, $50$, $71$, and $45$ frames in each sequence, respectively. For the ISBI dataset, The U373 comprises glioblastoma-astrocytoma U373 cells in a $115$ frames sequence of size $520\times 696$. Dataset GOWT1 documents GFP-GOWT1 mouse stem cells into $92$ frames with a size of $1024\times 1024$ per sequence. Huh7 is a dataset of human hepatocarcinoma-derived cells in a sequence of $30$ frames with a $1024\times 1024$ size.
@@
For the training process, we use $85$ training sequences in DeepCell and sequence $02$ in the ISBI dataset with only tracking (TRA) GT, from each dataset with cropping sequence $T_{s}=24$ frames for each iteration using sliding window length $l_{\text{slide}}=8$. We utilize pre-trained checkpoints provided by the authors of CoTracker [32], which has been trained exclusively on synthetic data with $3$ channels, so we duplicate grayscale images channel $3$ times. We use an end-to-end trained $2$-dimensional CNN architecture as the encoder to extract tracking features $F$. Specifically, it integrates bilinear interpolation and utilizes it to extract $S=4$ different scaled image features with $128$-channel output features concatenated to optimized dynamic feature extraction (Sec. 3.2). While training the transformer, we follow the pre-trained checkpoint that downsamples the resolution of frames to $384\times 512$ and choose $S=4$, $\Delta=3$ to calculate 4D multi-scaled correlation volumes (Sec. 3.2). More details can be found in the supplementary.
As for the testing (evaluation) process, we first select query points $Q$ for each cell that appeared on the first frame for each sequence and start the process of window rolling. The evaluation sequence $S_{\text{infer}}$, which is the subsequence $01$ in the dataset, follows the native resolution ($384\times 512$), which the dataset is trained on. We set $l_{\text{win}}=100$, which is mentioned in Sec. 3.3. We use the $\mathrm{TRA}$ evaluation criteria provided by ISBI [43, 62], which lay in the range $[0,1]$ that a higher score corresponds with a better performance and is calculated by $\mathrm{TRA}=1-\frac{\mathrm{min}(AOGM,AOGM_{0})}{AOGM_{0}}$, where $AOGM$ is the measure of acyclic oriented graphs matching [44], penalizing the number of transformations needed to transform the predicted tracking graph into the ground-truth tracking graph, e.g., merged cells ($\mathrm{EdgeSplit}$), missing links ($\mathrm{EdgeAdd}$), wrong links ($\mathrm{EdgeDelete}$), links with wrong semantics ($\mathrm{EdgeSemantic}$), false positives ($\mathrm{FP}$), and false negatives ($\mathrm{FN}$), and $AOGM_{0}$ is the $AOGM$ value for empty tracking results.
@@
Figure 5: Qualitative Result. Each point represents a cell with distinct colors, indicating individual cell trajectories across time points.
@@ 4.2 Results and Analysis
Tab. 1 presents the quantitative performance comparison of our method, CAP, and baseline methods, GNN, EmbedTrack, and KIT-GE on the ISBI dataset (U373, GOWT1, and Huh7) and DeepCell dataset (HeLa, PC-3, 3T3, and RAW264). Since $\mathrm{TRA}$ is derived from metrics $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, $\mathrm{FP}$, and $\mathrm{FN}$, we primarily focus on the $\mathrm{TRA}$ score for the model performance and $\mathrm{IT}$ for efficiency.
@@
U373 dataset with a $115$-frame sequence not only does CAP outperform all GNN, KIT-GE, and EmbedTrack significantly, achieving a perfect score of 00 for $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, and $\mathrm{FN}$ with the highest $\mathrm{TRA}$ score of $0.985$, but also it has the smallest inference time of $2.6$s. In contrast, GNN, EmbedTrack, and KIT-GE exhibit higher error counts and inference time. The $\mathrm{TRA}$ of datasets HeLa, PC-3, and RAW264, $0.926$, $0.952$, and $0.864$, show that The CAP not only significantly leads other methods in terms of the $\mathrm{TRA}$ values but also utilizes markedly less inference time compared to other approaches. CAP still demonstrates comparable performance with the $\mathrm{TRA}$ score of $0.960$, $0.966$, and $0.832$ across previous works in datasets GOWT1, Huh7, and 3T3. The trade-off remains advantageous given that CAP directly eliminates the stages of segmentation or detection, thereby significantly reducing the volume of data labels required and inference time ($7.1$s, $3.2$s, and $2.2$s) by approximately $5.5$ to $39$ times, while it has a slightly lower $\mathrm{TRA}$ measure than the greatest performance in existing methods.
@@
As for the extreme instability of EmbedTrack method on Huh7 with $\mathrm{TRA}=0.187$, our analysis suggests that the observed results stem from the method’s excessive reliance on the segmentation (SEG) stage and the usage of refined masks during the training of the segmentation model, while Huh7 contains only SEG golden truth (GT) masks without the more refined SEG silver truth (ST). More details about the difference between tracking GT mask, SEG GT, and SEG ST is illustrated in the supplementary. This further substantiates an end-to-end one-stage cell tracking framework, which bypasses the segmentation or detection stage, enhances model stability, and raises the overall performance limits achievable by the model.
@@
Overall, the CAP method demonstrates robust performance and constant high efficiency on inference and label usage across all datasets, consistently achieving high $\mathrm{TRA}$ scores and the lowest $\mathrm{IT}$s. In other words, these results suggest that CAP is a highly effective, efficient, and reliable method for cell tracking field in a practical biomedical context. Fig. 5 demonstrates a great qualitative result (visualization) of a sequence of time-lapse microscopy images obtained by CAP cell tracking framework.
@@ 4.3 Ablation Study
By replacing the cross-trajectory attention layers with only time attention layers to keep the model size, Tab. 3 shows model performance declined by $0.015$ to $0.055$ across datasets. Thus, the mechanism of tracking cells jointly plays an important role in CAP. More ablations can be found in the supplementary.
@@ 5 Conclusion and Future Work
In this paper, we introduce the one-stage CAP (Cell as Point) framework, which eliminates the need for a separate segmentation or detection stage—a significant innovation in cell tracking. This approach substantially improves efficiency by reducing inference time by a factor of $10$ to $55$, while also minimizing the need for high-quality data annotation. CAP effectively captures complex cell states and tracks cells in a single step, all while maintaining high efficiency. Additionally, it addresses common challenges in cell tracking, such as data imbalance and long sequence inference. In conclusion, CAP not only offers considerable practical value in real-world applications due to its efficiency but also pioneers a novel approach in the field of cell tracking.
@@
Appendix D: Additional Qualitative Results
- Provides more qualitative results (visualization) on detailed cell division events.
@@ B.2 DeepCell Dataset
Figure 8: Qualitative Result of Cell Division Details. Cell #1 is the mother cell, and cells #2 and #3 are daughter cells. (a) and (b) are two of the cell division events that occur in datasets HeLa and 3T3, respectively.
@@ C.2 Evaluation Metrics
Appendix D Additional Qualitative Results
In the main manuscript, we demonstrate the general qualitative result of CAP cell tracking performance. To show more details, i.e., the process of cell division, Fig. 8 demonstrates that CAP is capable of capturing cell division events and continuously tracking the new cells in different datasets in the long term.

### v2

@@ 3.3 Rolling-as-Window (RAW) Inference
Table 1: Comparison Experiments with Baselines. $\mathrm{ES}$, $\mathrm{EA}$, $\mathrm{ED}$, and $\mathrm{ESM}$ are abbreviations of $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, and $\mathrm{EdgeSemantic}$, respectively. Our framework CAP is TRA only, while previous frameworks GNN, EmbedTrack, and KIT-GE are SEG+TRA illustrated in Fig. 1 (a). Due to space limitations, KIT-GE results on DeepCell datasets are included in the supplementary. Average inference time/seconds ($\mathrm{IT}$) and performance ($\mathrm{TRA}$) are visualized in Fig. 1 (b).
In cell tracking, the inference stage is crucial for producing the final tracking mask and establishing cell lineage relationships by analyzing predicted cell locations and division events. Existing works [68, 64, 78, 49, 8] often customize inference and post-processing stages with different configurations for each dataset, adapting to variations in sequence length. This data-specific tailoring limits their generalizability and broader applicability to other datasets. In contrast, our proposed Rolling-As-Window (RAW) inference approach in CAP simplifies and unifies these processes while maintaining effectiveness across varying sequence lengths, enabling high-quality inference regardless of computational constraints. Specifically, a consistent and high-quality inference result is achieved through a rolling loop, which sequentially extracts yet-to-be-inferred clips from the full sequence, processing them within a defined window for inference.
4 Experiments and Results
We conducted experiments to demonstrate the high efficacy and efficiency of CAP, our end-to-end, one-stage framework for cell tracking. Despite requiring minimal inference time and eliminating the need for additional segmentation masks, CAP achieves promising performance. Additionally, we perform comprehensive ablation studies to assess the effectiveness of individual CAP components.
@@ 4.1 Experimental Setup
Figure 4: Cell Trajectories. Each point represents a cell with distinct colors, indicating individual cell trajectories across time points.
Figure 5: Cell Divisions. (a) and (b) illustrate cell division events accurately detected by the CAP framework in the HeLa and 3T3 datasets, respectively.
@@
We evaluate our proposed framework on DynamicNuclearNet Tracking in DeepCell [54] public database, HeLa, PC-3, 3T3, and RAW264 datasets divided by specimen, and ISBI cell tracking challenge [52, 74] public datasets, PhC-C2DH-U373 (U373), Fluo-N2DH-GOWT1 (GOWT1), and Fluo-C2DL-Huh7 (Huh7). These temporal data used for cell tracking, shot by different devices in various durations on diverse specimens, exhibit diverse morphologies and sequence characteristics. DeepCell contains $96$ sequences in total: $45$ for HeLa, $7$ for PC-3, $10$ for 3T3, and $34$ for RAW264, with $42$, $50$, $71$, and $45$ frames in each sequence, respectively. For the ISBI dataset, the U373 comprises glioblastoma-astrocytoma U373 cells in a $115$-frame sequence of size $520\times 696$. Dataset GOWT1 documents GFP-GOWT1 mouse stem cells into $92$ frames with a size of $1024\times 1024$ per sequence. Huh7 is a dataset of human hepatocarcinoma-derived cells in a sequence of $30$ frames with a $1024\times 1024$ size.
@@
For the training process, we utilize $85$ training sequences from the DeepCell dataset and sequences $02$ from the ISBI dataset, using only tracking (TRA) ground truth (GT). During each iteration, we crop sequences of $T_{s}=24$ frames with a sliding window length of $l_{\text{slide}}=8$. We use pre-trained checkpoints from CoTracker [40], which was trained solely on synthetic RGB data. Since our input images are grayscale, we replicate the single channel three times to match the required RGB format. We use an end-to-end trained $2$-dimensional CNN architecture as the encoder to extract tracking features $F$. Specifically, it integrates bilinear interpolation and utilizes it to extract $S=4$ different scaled image features with $128$-channel output features concatenated to optimized dynamic feature extraction (Sec. 3.2). While training the transformer, we follow the pre-trained checkpoint that downsamples the resolution of frames to $384\times 512$ and choose $S=4$, $\Delta=3$ to calculate 4D multi-scaled correlation volumes (Sec. 3.2). More details can be found in the supplementary.
As for the testing (evaluation) process, we preprocess the first frame for each sequence using [64] to select query points $Q$ and start the process of window rolling. The evaluation sequence $S_{\text{infer}}$ follows the native resolution ($384\times 512$), which the dataset is trained on. We set $l_{\text{win}}=100$, which is mentioned in Sec. 3.3. We use the $\mathrm{TRA}$ evaluation criteria provided by ISBI [52, 74] and CTMC [4], which lay in the range $[0,1]$ that a higher score corresponds with a better performance and is calculated by $\mathrm{TRA}=1-\frac{\mathrm{min}(AOGM,AOGM_{0})}{AOGM_{0}}$, where $AOGM$ is the measure of acyclic oriented graphs matching [53], penalizing the number of transformations needed to transform the predicted tracking graph into the ground-truth tracking graph, e.g., merged cells ($\mathrm{EdgeSplit}$), missing links ($\mathrm{EdgeAdd}$), wrong links ($\mathrm{EdgeDelete}$), links with wrong semantics ($\mathrm{EdgeSemantic}$), false positives ($\mathrm{FP}$), and false negatives ($\mathrm{FN}$), and $AOGM_{0}$ is the $AOGM$ value for empty tracking results.
@@ 4.2 Results and Analysis
Tab. 1 presents the quantitative performance comparison of our method, CAP, and baseline methods, GNN, EmbedTrack, and KIT-GE on the ISBI dataset (U373, GOWT1, and Huh7) and DeepCell dataset (HeLa, PC-3, 3T3, and RAW264). Since $\mathrm{TRA}$ is derived from metrics $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, $\mathrm{FP}$, and $\mathrm{FN}$, we primarily focus on the $\mathrm{TRA}$ score for the model performance and $\mathrm{IT}$ for efficiency (highlighted in gray).
@@
U373 dataset with a $115$-frame sequence not only does CAP outperform all GNN, KIT-GE, and EmbedTrack significantly, achieving a perfect score of 00 for $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, and $\mathrm{FN}$ with the highest $\mathrm{TRA}$ score of $0.985$, but also it has the smallest inference time of $2.6$s. In contrast, GNN, EmbedTrack, and KIT-GE exhibit higher error counts and inference time. The $\mathrm{TRA}$ of datasets HeLa, PC-3, and RAW264, $0.926$, $0.952$, and $0.864$, show that The CAP not only significantly leads other methods in terms of the $\mathrm{TRA}$ values but also utilizes markedly less inference time compared to other approaches. CAP still demonstrates comparable performance with the $\mathrm{TRA}$ score of $0.960$, $0.966$, and $0.832$ across previous works in datasets GOWT1, Huh7, and 3T3. The trade-off remains advantageous given that CAP directly eliminates the stages of segmentation or detection, thereby significantly reducing the volume of data labels required and inference time ($7.1$s, $3.2$s, and $2.2$s) by approximately $5.5$ to $39$ times, while it has a slightly lower $\mathrm{TRA}$ measure than the existing method. Fig. 4 highlights CAP’s robust cell trajectory tracking in time-lapse microscopy, while Fig. 5 demonstrates its strong capability in detecting cell divisions, accurately tracking mother and daughter cells across frames.
@@
As for the extreme instability of EmbedTrack method on Huh7 with $\mathrm{TRA}=0.187$, our analysis suggests that the observed results stem from the method’s excessive reliance on the segmentation (SEG) stage and the usage of refined masks during the training of the segmentation model, while Huh7 contains only SEG golden truth (GT) masks without the more refined SEG silver truth (ST). The supplementary Appendix B illustrates more details about the difference between tracking GT mask, SEG GT, and SEG ST. This further substantiates an end-to-end one-stage cell tracking framework, which bypasses the segmentation or detection stage, enhances model stability, and raises the overall performance limits achievable by the model.
@@
Overall, the CAP method demonstrates robust performance and constant high efficiency on inference and label usage across all datasets, consistently achieving high $\mathrm{TRA}$ scores and the lowest $\mathrm{IT}$s. Thus, CAP is a highly effective, efficient, and reliable method for cell tracking in a practical biomedical context.
@@ 4.3 Ablation Study
By replacing the cross-trajectory attention layers with only time attention layers to keep the model size, Tab. 3 shows model performance declined by $0.015$ to $0.055$ across datasets. Thus, cell joint tracking plays an important role in CAP.
@@
Additional ablations on DeepCell datasets are included in the supplementary.
@@ 5 Conclusion and Future Work
In this paper, we introduce the one-stage CAP (Cell as Point) framework, which eliminates the need for a separate segmentation or detection stage—a significant innovation in cell tracking. This approach substantially improves efficiency by reducing inference time by a factor of $10$ to $55$ and minimizing the need for high-quality data annotation. CAP effectively captures complex cell states and tracks cells in a single step, all while maintaining high efficiency. Additionally, it addresses common challenges in cell sequence data, such as cell event imbalance and long sequence inference. In conclusion, CAP not only offers considerable practical value in real-world applications due to its efficiency but also pioneers a novel approach in the field of cell tracking.
@@
Appendix D: Additional Baseline Results
- Provides KIT-GE baseline method results on the DeepCell dataset as complementary for Tab. 1 in Sec. 4.2.
@@ B.2 DeepCell Dataset
Metrics HeLa PC-3 3T3 RAW264
KIT-GE CAP KIT-GE CAP KIT-GE CAP KIT-GE CAP
$\mathrm{ES}$downarrow 6 3 29 1 128 107 37 67
$\mathrm{EA}$downarrow 57 82 56 11 783 949 585 1098
$\mathrm{ED}$downarrow 1 1 0 0 14 2 106 12
$\mathrm{ESM}$downarrow 17 0 0 0 10 4 24 17
$\mathrm{FP}$downarrow 70 60 13 35 256 250 718 647
$\mathrm{FN}$downarrow 67 70 1 7 698 709 603 898
$\mathrm{TRA}$uparrow 0.923 0.926 0.937 0.952 0.820 0.832 0.876 0.864
$\mathrm{IT}$downarrow 14.4 1.3 9.1 1.1 19.1 2.2 96.7 2.8
Table 6: Comparison Experiment with Baseline. Here is a complementary for Tab. 1 in Sec. 4.2.
@@ C.2 Evaluation Metrics
Appendix D Additional Baseline Results
Tab. 6 provides additional baseline results as a complementary analysis to Tab. 1 in Sec. 4.2., further demonstrating the effectiveness of our CAP framework. It shows the performance comparison of our proposed CAP framework against KIT-GE across four datasets, HeLa, PC-3, 3T3, and RAW264, in the DeepCell. CAP consistently achieves higher TRA (tracking accuracy) scores than KIT-GE in HeLa, PC-3, and 3T3 datasets, demonstrating its superior tracking capability. Furthermore, CAP significantly reduces IT (inference time) across all datasets, being 10×10$\times$10 × to $\times8$ faster than KIT-GE, highlighting its efficiency. Overall, CAP delivers superior tracking accuracy with significantly lower inference time, reinforcing its effectiveness as a one-stage tracking solution compared to the multi-stage KIT-GE framework.